OpenStudy (anonymous):

I need help. Can you write a rule for the linear function in the graph? A. y = -4x + 13 B. y = -4x – 13 C. y = 4x – 13 D. y = 1/4x – 13

OpenStudy (anonymous):

@Nnesha ?

OpenStudy (aum):

Is there a graph that goes with the first question?

OpenStudy (anonymous):

Sorry @aum, here is the graph.

OpenStudy (anonymous):

I uploaded the graph, @aum

OpenStudy (phi):

they give you two points on the line (3,-1) and (4,3) do you know how to find the slope between those points?

OpenStudy (anonymous):

no, sorry =(

OpenStudy (phi):

do you know that they are (x,y) pairs? can you pick out the y values for those two points ?

OpenStudy (anonymous):

3 and -1???

OpenStudy (phi):

yes. the first step is find the "change in y" which means subtract can you do that ?

OpenStudy (anonymous):

Subtract 3 and -1?

OpenStudy (phi):

yes

OpenStudy (anonymous):

4?

OpenStudy (phi):

3 - -1 = 3+1 = 4 yes now find the x values of the two points and find the "change in x"

OpenStudy (anonymous):

4-3=1?

OpenStudy (aum):

Sorry the notifications don't work too well. Just got the notification 25 minutes later.

OpenStudy (phi):

yes finally, slope is change in y divided by change in x

OpenStudy (anonymous):

lol its okay @aum =)

OpenStudy (anonymous):

0.25?

OpenStudy (phi):

what did you get for change in y ?

OpenStudy (anonymous):

4

OpenStudy (phi):

and 1 for change in x now do "change in y" over "change in x"

OpenStudy (anonymous):

4/1

OpenStudy (phi):

and that simplifies to just 4

OpenStudy (phi):

you found the slope is 4 now we use y = m x + b that is the formula of a line. m is the slope. We could use the info to also find b. but so far, we know this: y= 4 x + b so your choice better have a 4x in it.

OpenStudy (anonymous):

okay cool!

OpenStudy (phi):

only 1 choice works. but if we wanted to find the "b" in y = 4 x + b use one of the points, say (3,-1) which says when x is 3, y is -1. use those in the equation: -1 = 4*3 + b 4*3 is 12 -1 = 12 + b now add -12 to both sides -1 - 12 = 12 - 12 + b simplify to get -13 = b and the full equation is y = 4x -13

OpenStudy (anonymous):

WOW! Thank you for your brain! =)